# SciCoQA: Quality Assurance for Scientific Paper–Code Alignment

### v1

@@ 3.2 Data Analysis
Fig. 3 analyzes several dimensions of the SciCoQA dataset. While our real-world data is mostly from computer science (specifically from AI and its subdomains such as Machine Learning, Computer Vision, and Natural Language Processing), the synthetic data contains papers and code from Electrical Engineering and System Science, Physics, Statistics, Quantitative Biology, and Mathematics, making the data diverse across computational sciences.
We further analyze the discrepancy type (as defined in §3), by annotating each discrepancy in the real-world set as Difference, Paper Omission, or Code Omission. For the synthetic data, we provide discrepancy type definitions to GPT-5 in the prompt and let it generate the label along with the discrepancy. We find that $53\%$ of discrepancies in the real data are Differences, which is also the majority class in the synthetic data, accounting for $80\%$ of the data.
Beyond the discrepancy type, we analyze the affected component within the research pipeline. We introduce a taxonomy of six discrepancy categories: Algorithm: changes to step order, operations, or core logic; Model: architectural or initialization changes; Loss: alterations to loss definitions or terms; Evaluation: modifications to evaluation logic, metrics, or scripts; Data: dataset usage, preprocessing, augmentation, or filtering; and Training: changes to the learning process, schedule, or optimization. We apply this taxonomy to real-world data and the CS subset of synthetic data, excluding other scientific domains (e.g., Physics) where these ML-specific concepts may not always apply cleanly. Similar to the discrepancy type annotation, we manually label the real instances and automatically generate labels for the synthetic ones. In the real data, discrepancies in the Algorithm and Loss dominate, with $26\%$ each, whereas the synthetic data also has $26\%$ of discrepancies in the Algorithm, and Model discrepancies are the second-highest category, at $20\%$.
We further visualize the number of tokens in the prompt. Papers have a median of $14,350$, codebases $38,978$, and the combined prompt $57,008$ tokens. Further, $70$ out of $266$ papers have more than $100k$ tokens combined with their codebase, making SciCoQA a challenging task to measure the models’ long-context abilities. We analyze the data by programming languages (§E) and the synthetic code (§D) further in the appendix.
@@ 4 Experiments
Given a paper and its code, we prompt the model to generate a list of discrepancies between the two (prompt in §B.6). We then parse the model output into individual discrepancies. The generation prompt contains the same instructions as those used to construct the ground-truth discrepancies. For implementation details, we refer to §C. We further ablate the discrepancy prediction experiment by providing only the code as input. This allows us to quantify the contribution of the paper, distinguishing between discrepancies that require cross-modal reasoning and those that can be inferred from the code (e.g., through comments or readmes), or the model’s parametric knowledge of the paper.
@@
We employ LLM-as-a-Judge (Zheng et al., 2023) to evaluate whether the predicted discrepancy matches the reference. Inspired by Wei et al. (2025), we use a reasoning model, specifically GPT-OSS 20B (OpenAI, 2025b) (prompt in §B.7), because it is open-weight, enabling reproducibility, and due to its favorable speed-performance trade-off. To verify the evaluation setup, we annotate the predictions from $20$ discrepancies, where $10$ originate from GitHub and $10$ from reproducibility papers. To streamline our evaluation and spend most time on annotating relevant generations, we compute the top three most similar predicted discrepancies per model using EmbeddingGemma (Vera et al., 2025). We then manually assess whether each prediction matches the reference discrepancy. In total, this yields $1,039$ annotations, of which $143$ predictions match the reference. On this data, the LLM judge achieves an F1 score of $87.5\pm 1.1$, showing strong alignment with our annotations. [^3] We observe a slight variance in predictions despite a fixed seed in vLLM, we therefore report the mean and standard deviation over five runs. While the evaluation is a binary classification (reference and predicted discrepancies match or do not match), we refer to the performance as recall, as the predictions may contain unannotated, but valid, discrepancies, and we only evaluate whether the annotated ones are detected.
@@
We evaluate several state-of-the-art model families, including commercial and open models, as well as reasoning, instruction-tuned, and code-specific variants, specifically: GPT-5, Gemini 2.5, GPT-OSS, Qwen3, DeepSeek R1, Nemotron, and Mistral (exact model variants in Table 3).
Figure 4: Results of the top 8 best performing models (sorted by average recall on the real and synthetic data) on the discrepancy dataset by different analyses. From left to right, we analyze the origin of the discrepancy, the type of discrepancy, the number of tokens in the input prompt, and the publication year of the paper.
@@ 5 Results
Fig. 4 reports the recall of the top-performing models (full results in Table 5). Overall, GPT-5 performs best, achieving a recall of $45.7\%$, but leaving considerable room for improvement. On the synthetic data, $71.3\%$ are detected by the model. Crucially, we observe a strong correlation ($r=0.94$) between recall on real and synthetic data (see §F.1), validating the synthetic subset as a reliable proxy for real-world discrepancies. While general scaling laws can be observed (i.e., models trained with more compute and data perform better), GPT-OSS 20B is an exception on our task, outperforming its 120B variant. We attribute this, on the one hand, to the self-preference bias of the judge, rating its own outputs better (Panickssery et al., 2024), and on the other hand, to the model’s verbosity. GPT-OSS 20B makes, on average, $5.1$ predictions per paper, while the 120B model only makes $4.5$, giving the smaller variant a higher chance to match the discrepancy at the expense of precision. Additionally, we find that GPT-5 Mini outperforms GPT-5 Codex, despite the latter being the larger model (Codex is based on GPT-5). While Codex is generally superior in code generation, for SciCoQA code and natural language understanding are both crucial, and we conjecture that the general instruction-following and reasoning abilities of GPT-5 and GPT-5 Mini are more helpful than specialized coding knowledge.
@@
Analyzing detection rates reveals a clear hierarchy: Code Omissions are the easiest to detect, followed by Differences, while Paper Omissions are the most challenging. This order also explains the performance gap between data sources. Discrepancies from GitHub are easier to detect because they primarily consist of Differences ($68\%$) and Code Omissions ($21.6\%$). In contrast, discrepancies from reproducibility papers are more challenging as they are often dominated by Paper Omissions ($50\%$). This likely stems from the asymmetry between modalities: the paper acts as a specification, where all described components must exist in the code. Code Omissions and Differences benefit from this explicit grounding. Conversely, the code contains many implementation details not required in the paper; thus, detecting Paper Omissions requires the harder task of distinguishing deviations from permissible engineering artifacts without a textual reference.
@@
We further analyze the performance by the number of input tokens. We split the dataset by prompt length into five approximately equally sized buckets. We find a consistent pattern where the longer the input, the lower the performance becomes, as similarly observed in various studies (Hsieh et al., 2024; Zhang et al., 2024). This sensitivity to input length also suggests an explanation for the performance between domains in the synthetic data: models struggle more with CS papers, which is likely driven by the larger repository size of CS vs non-CS repositories (Median: $50k$ vs $28k$ tokens). This increased volume creates a larger search space, making the task more challenging. While many long context evaluations rely on artificially constructed “needles” that need to be retrieved from a long context, i.e., the haystack (Kamradt, 2023), SciCoQA can offer a natural evaluation setup similar to Liu et al. (2024a) for long-context evaluation of LLMs. [^4] The synthetic data generation can be used to inject discrepancies at controlled places in the codebase.
@@
Lastly, we split the data by publication year. Among the top models, the one with the most recent knowledge cutoff is the Gemini 2.5 family, which was trained using data up to January 2025 (other models have cutoffs in 2024). Therefore, the 2025 split of our data can be considered not part of the pre-training data. Most models (except the Gemini 2.5 Flash) perform the worst on the most recent data. This demonstrates that models benefit from having the paper and code included in their pre-training data, aligning with previous findings on data contamination (OpenAI, 2023; Zhou et al., 2023). This also emphasizes the importance of our synthetic data generation pipeline, which enables us to continuously update our dataset with uncontaminated data that is not part of the pre-training data of future models.
@@
Beyond the top-performing models, we find the other open-weight models (Qwen3, Nemotron, DeepSeek R1, Mistral) severely limited. Nemotron 49B and Qwen3 Coder 30B perform best, but only achieve an average recall on the combined data of $23.6\%$ and $23.1\%$, respectively.
In summary, SciCoQA remains a significant challenge for state-of-the-art LLMs. We find the recall of all models to be limited; the best-performing model can only detect $45.7\%$ of all discrepancies in the real-world data. Consequently, these systems cannot yet be relied upon to ensure the faithfulness of scientific publications with their codebases.
@@ 5.1 Code Only Ablation
To test the multimodality of our data, we remove the paper from the input, leaving only the code. This makes it more difficult to identify discrepancies, but not impossible since the repositories also frequently contain abstracts, summaries, or details of the paper in their readme files. Furthermore, scientific papers are typically part of the pre-training data; therefore, while the model does not directly receive the paper as input, it may be able to recall it from its pre-training. We show the results of the top models in this experiment in Fig. 5 (full results in Table 6). When removing the paper from the input, all models perform worse, confirming that the paper is necessary to perform the task. We observe an average drop of $19.4$ percentage points, a relative drop of $53.4\%$ on the real data. For the synthetic data, the drop is less pronounced; on average, the models’ performance drops by $15.1$ percentage points (relative drop of $30.3\%$).
@@
False Positives 88 22 2121
OCR Error 11 11 11
Paper Ambiguity 22 11 0
Paper Misunderstanding 22 0 66
Code Misunderstanding 0 0 $12$
3rd Party Code 22 0 11
True Positives $35$ $26$ $43$
Unique $15$ $10$ $24$
Minor 66 55 $12$
Total $43$ $28$ $64$
Precision $81.4$ $92.9$ $67.2$
Table 1: Analysis of unlabeled discrepancy predictions of GPT-5, Gemini 2.5 Pro, and GPT-OSS 20B of 12 NLP papers from the real subset of SciCoQA.
5.2 Validation of Unlabeled Discrepancies
To better understand the precision of the models, we analyze GPT-5’s, Gemini 2.5 Pro’s, and GPT-OSS 20B’s generations. We select the predictions for $12$ NLP papers that were not matched to a ground truth, yielding a total of $135$ discrepancies (results in Table 1). We select NLP papers to leverage our own domain expertise, as validating discrepancies is challenging and requires a detailed understanding of the paper and its implementation.
Overall, we find that Gemini 2.5 Pro does not make significant errors, achieving a precision of $92.9\%$. Its two errors were due to preprocessing (an omitted term in a formula during OCR) and an ambiguity within the paper, a nuance even noted in the model’s generation. GPT-5 followed ($81.4\%$), primarily failing due to paper ambiguities or misinterpretations, or incorrect assumptions about third-party library functions. Finally, GPT-OSS 20B struggles most ($67.2\%$), frequently misunderstanding the code logic (e.g., it fails to align variables that are named differently in the paper and code or misinterprets conditional execution paths).
We further construct a pooled ground truth for the $12$ analyzed papers. We aggregate the discrepancies originally annotated in SciCoQA with the verified ones identified by the models (excluding minor ones, which are mostly discrepancies related to mismatching settings that can typically be resolved via configuration or the CLI). This results in a comprehensive set of $67$ distinct paper-code discrepancies (results in Table 2). We find that the number of discrepancies uniquely identified by the models is high (see also Table 1), and recall remains low ($43$-$60\%$), confirming our main experiment’s findings that model recall is limited. GPT-5 achieves the best precision-recall tradeoff, obtaining an F1 score of $67.3\%$.
Model GPT-5 Gemini GPT-OSS
True Positives $38$ $29$ $40$
False Positives 88 22 2020
Precision $82.6$ $93.5$ $66.7$
Recall $56.7$ $43.3$ $59.7$
F1 $67.3$ $59.2$ $63.0$
Table 2: Performance of GPT-5, Gemini 2.5 Pro, and GPT-OSS 20B on 67 pooled discrepancies from the annotation of unlabeled predictions and discrepancies from SciCoQA from 12 NLP papers.
@@ 6 Conclusion
We introduced SciCoQA, a dataset of 611 discrepancies designed to evaluate the alignment between a scientific paper and its code. Our analysis reveals a critical gap: while models like GPT-5 demonstrate high precision, they suffer from insufficient recall, detecting only $45.7\%$ of real-world discrepancies. We find that models particularly struggle with paper omissions, where code logic is absent from the paper text, and fail to maintain performance on recent, uncontaminated publications. Consequently, while current LLMs show promise as assistants, they cannot yet serve as autonomous arbiters of scientific validity. As research scales toward “AI Scientists,” SciCoQA provides the essential ground truth to ensure these agents remain faithful to the scientific method. To further advance this field, we encourage the broader adoption of reproducibility tracks to systematically capture these discrepancies and expand future data collection.
Limitations

### v2

@@ 3.2 Data Analysis
Fig. 3 analyzes several dimensions of SciCoQA. While our real-world data is mostly from computer science (specifically from AI and its subdomains such as ML, CV and NLP), the synthetic data contains papers and code from Electrical Engineering and System Science, Physics, Statistics, Quantitative Biology, and Mathematics, making the data diverse across computational sciences.
We further analyze the discrepancy type (as defined in §3), by annotating each discrepancy in the real set as Difference, Paper Omission, or Code Omission. For the synthetic data, we provide type definitions in the prompt and generate the label along with the discrepancy. We find that $55\%$ of discrepancies in the real data are Differences, which is also the majority class in the synthetic data, accounting for $78\%$ of the data.
Beyond the discrepancy type, we analyze the affected component within the research pipeline. We introduce a taxonomy of six discrepancy categories: Algorithm: changes to step order, operations, or core logic; Model: architectural or initialization changes; Loss: alterations to loss definitions or terms; Evaluation: modifications to evaluation logic, metrics, or scripts; Data: dataset usage, preprocessing, augmentation, or filtering; and Training: changes to the learning process, schedule, or optimization. We apply this taxonomy to real-world data and the CS subset of synthetic data, excluding the other domains where these ML-specific concepts may not always apply cleanly. Similar to the discrepancy type annotation, we manually label the real instances and automatically generate labels for the synthetic ones. In the real data, discrepancies in the Algorithm and Loss dominate ($25\%$ and $24\%$) whereas the synthetic data has $26\%$ of discrepancies in the Algorithm, and Model discrepancies are the second-highest category, at $21\%$.
Finally, we visualize the number of tokens in the prompt. Papers have a median of 14,20014{,}200, codebases 39,27239{,}272, and the combined prompt 56,90356{,}903 tokens. Further, $73$ out of $276$ papers have more than $100k$ tokens combined with their codebase, making SciCoQA a challenging task to measure the models’ long-context abilities. We analyze the data by programming languages (§B.3) and the synthetic code (§B.2) further in the appendix.
@@ 4 Experiments
Given a paper and its code, we prompt the model to generate a list of discrepancies between the two (prompt in §F.6). We then parse the model output into individual discrepancies. The generation prompt contains the same instructions as those used to construct the ground-truth discrepancies. For implementation details, we refer to §A. We further ablate the discrepancy prediction experiment by providing only the code as input. This allows us to quantify the contribution of the paper, distinguishing between discrepancies that require cross-modal reasoning and those that can be inferred from the code (e.g., through comments or readmes), or the model’s parametric knowledge of the paper.
@@
We employ LLM-as-a-Judge (Zheng et al., 2023) to evaluate whether the predicted discrepancy matches the reference. Inspired by Wei et al. (2025a), we use a reasoning model, specifically GPT-OSS 20B (OpenAI, 2025b) (prompt in §F.7), because it is open-weight, enabling reproducibility, and its favorable speed-performance trade-off. For the real data where two ground truth descriptions exists (generated by Gemini 3.1 Pro and GPT-5), we consider matches against either as correct. To verify the evaluation setup, we annotate the predictions from $20$ discrepancies, where $10$ originate from GitHub and $10$ from reproducibility papers. To streamline our evaluation, we compute the top three most similar predicted discrepancies per model using EmbeddingGemma (Vera et al., 2025). We then manually assess whether each prediction matches the reference discrepancy. In total, this yields 1,0391{,}039 annotations, of which $143$ predictions match the reference. On this data, the LLM judge achieves an F1 score of $87.5\pm 1.1$, showing strong alignment with our annotations. [^3] We observe a slight variance in predictions despite a fixed seed in vLLM, we therefore report the mean and standard deviation over five runs. We further validate the setup by running Qwen3 32B as an alternative judge. We find Cohen’s Kappa between them at $\kappa=83.8$ and an F1 score of $91.9$. However, since the F1 score of Qwen3 32B with our human evaluation is at $70.0$, we retain GPT-OSS 20B as a judge. While the evaluation is a binary classification (reference and predicted discrepancies match or do not match), we refer to the performance as recall, as the predictions may contain unannotated, but valid, discrepancies, and we only evaluate whether the annotated ones are detected.
@@
We evaluate several state-of-the-art model families, including commercial and open models, as well as reasoning, instruction-tuned, and code-specific variants, specifically: GPT-5, Gemini, GPT-OSS, Qwen3, DeepSeek R1, Nemotron, and Mistral (exact model variants in Table 2).
Figure 4: Results of the top 8 best-performing models (sorted by average recall on the real and synthetic data) on the discrepancy dataset by different analyses. From left to right: we analyze the origin of the discrepancy, the type of discrepancy, the number of tokens in the prompt, and the publication year of the paper.
@@ 5 Results
Fig. 4 reports the recall of the top-performing models (full results in Table 9). Overall, Gemini 3.1 Pro and GPT-5 Mini perform best, achieving a recall of $46.7\%$, leaving considerable room for improvement. On the synthetic data, GPT-5 detects $70.0\%$ discrepancies. Crucially, we observe a strong correlation ($r=0.94$) between recall on real and synthetic data (see §C.4), validating the synthetic subset as a reliable proxy for real-world discrepancies. While general scaling laws can be observed (i.e., models trained with more compute and data perform better), GPT-5 Mini and GPT-OSS 20B are an exception in our task, outperforming their larger variants slightly. We attribute this, on the one hand, to the self-preference bias of the judge for GPT-OSS, rating its own outputs better (Panickssery et al., 2024), and on the other hand, to the model’s verbosity. GPT-5 Mini and GPT-OSS 20B on average make $5.3$ and $5.8$ predictions per paper, while GPT-5 and GPT-OSS 120B make only $4.6$ and $5.0$ predictions, giving the smaller variants a higher chance to match the discrepancy at the expense of precision. Additionally, we find GPT-5 Codex to be inferior, despite it being a larger model than GPT-5 Mini (Codex is based on GPT-5). While Codex is generally superior in code generation, for SciCoQA code and natural language understanding are both crucial, and we conjecture that the general instruction-following and reasoning abilities of GPT-5 and GPT-5 Mini are more helpful than specialized coding knowledge.
@@
Analyzing detection rates reveals a clear hierarchy: Code Omissions are the easiest to detect, followed by Differences, while Paper Omissions are the most challenging. This order also explains the performance gap between data sources. Discrepancies from GitHub are easier to detect because they primarily consist of Differences ($71.4\%$) and Code Omissions ($19.0\%$). In contrast, discrepancies from reproducibility papers are more challenging as they are often dominated by Paper Omissions ($50\%$). This likely stems from the asymmetry between modalities: the paper acts as a specification, where all described components must exist in the code. Code Omissions and Differences benefit from this explicit grounding. Conversely, the code contains many implementation details not required in the paper; thus, detecting Paper Omissions requires the harder task of distinguishing deviations from permissible engineering artifacts without a textual reference.
@@
We further analyze the performance by the number of input tokens.444§C.3 provides an analysis by relevant code file position. We split the dataset by prompt length into five approximately equally sized buckets. We find a consistent pattern where the longer the input, the lower the performance becomes, as similarly observed in various studies (Hsieh et al., 2024; Zhang et al., 2024; Levy et al., 2024). This sensitivity to input length also suggests an explanation for the performance between domains in the synthetic data: models struggle more with CS papers, which is likely driven by the larger repository size of CS vs non-CS repositories (median: $53k$ vs $31k$ tokens). This increased volume creates a larger search space, making the task more challenging. While many long context evaluations rely on artificially constructed “needles” that need to be retrieved from a long context, i.e., the haystack (Kamradt, 2023), SciCoQA can offer a natural evaluation setup similar to Liu et al. (2024a) for long-context evaluation of LLMs. [^5] The synthetic data generation can be used to inject discrepancies at controlled places in the codebase.
@@
Lastly, we split the data by publication year. Among the top models, the one with the most recent knowledge cutoff is the Gemini family, which was trained using data up to January 2025 (other models have cutoffs in 2024). Therefore, the 2025 split of our data can be considered not part of the pre-training data. Most models (except Gemini 2.5 Flash) perform worst on the most recent data. This demonstrates that models benefit from having the paper and code included in their pre-training data, aligning with previous findings on data contamination (OpenAI, 2023; Zhou et al., 2023). This further emphasizes the importance of our synthetic data generation pipeline, which enables us to continuously update our dataset with uncontaminated data that is not part of the pre-training data of future models.
@@
Beyond the top-performing models, we find the other open-weight models (Qwen3, Nemotron, DeepSeek R1, Mistral) severely limited. Nemotron 49B and Qwen3 Coder 30B perform best, but only achieve an average recall on the combined data of $23.9\%$ and $23.5\%$, respectively.
In summary, SciCoQA provides a significant challenge for state-of-the-art LLMs. We find the recall of all models to be limited; the best-performing models can only detect $46.7\%$ of all discrepancies in the real-world data. Consequently, these systems cannot yet be relied upon to ensure the faithfulness of scientific publications with their codebases.
@@ 5.1 Code Only Ablation
To test the multimodality of our data, we remove the paper from the input, leaving only the code. This makes it more difficult to identify discrepancies, but not impossible, since the repositories also frequently contain abstracts, summaries, or details of the paper in their readme files. Furthermore, scientific papers are typically part of the pre-training data; therefore, while the model does not directly receive the paper as input, it may be able to recall it from its pre-training. We show the results of the top models in this experiment in Fig. 5 (full results in Table 10). When removing the paper from the input, all models perform worse, confirming that the paper is necessary to perform the task. We observe an average drop of $19.2$ percentage points, a relative drop of $48.3\%$ on the real data. For the synthetic data, the drop is less pronounced; on average, the models’ performance drops by $16.3$ percentage points (relative drop of $30.8\%$).
5.2 Validation of Unlabeled Discrepancies
To better understand the precision of the models, we analyze GPT-5’s, Gemini 2.5 Pro’s, and GPT-OSS 20B’s generations. We select predictions of $20$ papers from NLP and CV that were not matched to a ground truth, yielding a total of $224$ evaluations (results in Table 3). We select these papers to leverage our own domain expertise, as validating discrepancies is challenging and requires a detailed understanding of the paper and its implementation.
Overall, we find that Gemini 2.5 Pro does not make significant errors, achieving a precision of $94.1\%$. Its three errors were due to preprocessing (an omitted term in a formula during OCR), an ambiguity within the paper, a nuance even noted in the model’s generation and one error due to the incorrect recall of another paper’s formula. GPT-5 follows ($85.7\%$), primarily failing due to paper ambiguities or misinterpretations, or incorrect assumptions about library functions. Finally, GPT-OSS 20B struggles most ($68.0\%$), frequently misunderstanding the code logic (e.g., it fails to align variables that are named differently in the paper and code or misinterprets conditional execution paths).
Taking inspiration from information retrieval benchmarks where incomplete relevant judgments are common (Buckley and Voorhees, 2004), we construct a pooled ground truth from the $20$ papers. We aggregate the discrepancies originally annotated in SciCoQA with the verified ones identified by the models (excluding minor ones, which are mostly discrepancies related to mismatching settings that can typically be resolved via configuration or the CLI). This results in a comprehensive set of $129$ distinct paper-code discrepancies (results in Table 1). We find that the number of discrepancies uniquely identified by the models is high (see also Table 3), and recall remains low ($41$-$56\%$), confirming our main experiment’s findings that model recall is limited. GPT-5 achieves the best precision-recall tradeoff, obtaining an F1 score of $64.7\%$. We evaluate all $22$ models against this pooled ground truth in §C.6, confirming that recall remains the primary bottleneck.
@@ 5.2 Validation of Unlabeled Discrepancies
True Positives $66$ $55$ $72$
False Positives 99 33 $31$
Precision $88.0$ $94.6$ $69.9$
Recall $51.2$ $41.1$ $55.8$
F1 $64.7$ $57.3$ $62.1$
Table 1: Performance of GPT-5, Gemini 2.5 Pro, and GPT-OSS 20B on $129$ pooled discrepancies from the annotation of unlabeled predictions and discrepancies from SciCoQA from $20$ NLP and CV papers.
@@ 6 Conclusion
We introduced SciCoQA, a dataset of $635$ discrepancies designed to evaluate the alignment between a scientific paper and its code. Our analysis reveals a critical gap: while models like GPT-5 and Gemini demonstrate high precision, they suffer from insufficient recall, detecting only $46.7\%$ of real-world discrepancies. We find that models particularly struggle with paper omissions, where code logic is absent from the paper text, and fail to maintain performance on recent, uncontaminated publications. Consequently, while current LLMs show promise as assistants, they cannot yet serve as autonomous arbiters of scientific validity. As research scales toward “AI Scientists,” SciCoQA provides the essential ground truth to ensure these agents remain faithful to the scientific method. To further advance this field, we encourage the broader adoption of reproducibility tracks to systematically capture these discrepancies and expand future data collection.
7 Limitations